### v1

# Mills’ constant is irrational

Abstract.
Let $\lfloor x\rfloor$ denote the integer part of $x$. In 1947, Mills constructed a real number $\xi$ greater than $1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is always a prime number for every positive integer $k$. We define Mills’ constant as the smallest real number $\xi$ satisfying this property. In this article, we determine that Mills’ constant is irrational. Moreover, we also obtain partial results on the transcendency of the constant.
Key words and phrases:
prime-representing constants, Mills’ constant, transcendental number theory
2020 Mathematics Subject Classification:
11J72, 11J81
1. Introduction
Let $\mathbb{N}$ denote the set of all positive integers. Let $\lfloor x\rfloor$ denote the integer part of $x\in\mathbb{R}$. A complex number $\alpha$ is transcendental if $P(\alpha)\neq 0$ for all polynomials $P\in\mathbb{Q}[x]\setminus\{0\}$.
Mills [12] constructed a constant $\xi>1$ such that
[…]
Such $\xi$ is not uniquely determined. In this article, we define Mills’ constant [^1] There are two types of definitions of Mills’ constant. Finch [7] implicitly introduced Mills’ constant as a real number $\xi$ satisfying (1.1). Our definition can be seen in [3]. as the smallest $\xi>1$ satisfying (1.1). It is well-defined from [13, Remark 4.2], which means that Mills’ constant exists. Assuming the Riemann hypothesis, Caldwell and Cheng [3] determined $6850$ digits of the decimal expansion of Mills’ constant which is approximately $1.3063778838\cdots$. It is an unsolved problem to determine whether the constant is rational or irrational. The goal of this article is to solve it.
Before stating the results, for every integer $c\geq 2$, we define $\xi_{c}$ as the smallest real number $\xi$ greater than $1$ such that $\lfloor\xi^{c^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$. By [11, Corollary 4] and [13, Remark 4.2], $\xi_{c}$ exists for every integer $c\geq 2$. A real algebraic integer greater than $1$ is a Pisot number if all of its conjugates over $\mathbb{Q}$ except for itself lie in the open unit disc. We obtain the following results.
Theorem 1.1.
For every integer $c\geq 4$, the constant $\xi_{c}$ is transcendental.
Theorem 1.2.
Either the constant $\xi_{3}$ is transcendental, or there exists a positive integer $m$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. In particular, $\xi_{3}$ is irrational.
For more generalization, let $(c_{k})_{k\in\mathbb{N}}$ be a real sequence with $c_{k}\geq 1$ for every $k\in\mathbb{N}$. We define $C_{k}=c_{1}\cdots c_{k}$ for every $k\in\mathbb{N}$ and
$\mathcal{W}(c_{k})=\{\xi>1\colon\lfloor\xi^{C_{k}}\rfloor\text{ is a prime number for every $k\in\mathbb{N}$}\}.$
An element $\xi\in\mathcal{W}(c_{k})$ is called a prime-representing constant (for short PRC). Wright [15] firstly investigated the geometric structure of $\mathcal{W}(c_{k})$. Especially, by his result [15, Theorem 5], $\mathcal{W}(c_{k})$ is uncountable under suitable conditions on $(c_{k})_{k=1}^{\infty}$. Therefore, there exist infinitely many transcendental numbers in $\mathcal{W}(c_{k})$, but we do not get any arithmetic properties of the smallest element in $\mathcal{W}(c_{k})$ from this result. Alkauskas and Dubickas [1, Theorem 1] firstly gave a result on the transcendency of specific PRCs. They succeeded in constructing a transcendental number in $\mathcal{W}(c_{k})$ if a sequence $(c_{k})_{k\in\mathbb{N}}$ of real numbers satisfying
[…]
They did not discuss whether the smallest element in $\mathcal{W}(c_{k})$ is transcendental or not. The author and Takeda [13, Theorem 1.3] showed that the smallest element in $\mathcal{W}(c_{k})$ exists and it is transcendental if a sequence $(c_{k})_{k\in\mathbb{N}}$ of integers satisfying
[…]
They also obtained a result on the algebraic independency of PRCs in [13]. The condition (iii) is essentially needed in these works, but we accomplish replacing (iii) with the condition $c_{k+1}\geq 4$ for infinitely many positive integers $k$ as follows.
Theorem 1.3.
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying
  1. (1)
    $c_{1}\geq 1$;
  2. (2)
    $c_{k+1}\geq 2$ for all $k\in\mathbb{N}$;
  3. (3)
    $c_{k+1}=b$ for infinitely many positive integers $k$.
Then the smallest real number in $\mathcal{W}(c_{k})$ exists. Let $\xi$ be the number, and the following properties are true:
  • if $b\geq 4$, then $\xi$ is transcendental;
  • if $b=3$, then either $\xi$ is transcendental, or $\xi^{C_{m}}$ is a Pisot number of degree $3$ for some $m\in\mathbb{N}$.
Theorem 1.3 immediately implies Theorems 1.1 and 1.2. We will prove Theorem 1.3 in Section 4. At the last of this section, let us ask several remaining questions.
Question 1.4.
Is $\xi_{2}$ rational or irrational?
Question 1.5.
Does there exist $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$?
If the answer to Question 1.4 is “no”, then $\xi_{3}$ is transcendental by Theorem 1.2. In Section 5, we will give assumptions leading to the transcendency of $\xi_{3}$. We will explain the main difficulties of Questions 1.4 and 1.5 in Remarks 3.6 and 4.4, respectively.
2. Strategy of proofs
In this section, we demonstrate a strategy of proofs. We first recall Mills’ construction, that is, we now construct $\xi>1$ such that $\lfloor\xi^{3^{k}}\rfloor$ is a prime number for every $k\in\mathbb{N}$.
Let $\mathcal{P}$ be the set of all prime numbers, and let $\theta=21/40$. We apply the following theorem given by Baker, Harman, and Pintz [2] [^2] Mills applied the result given by Ingham [8] which states that $p_{n+1}-p_{n}<Kp_{n}^{5/8}$ for some positive constant $K$. Thus, the construction of this section is modernly modified. .
Theorem 2.1.
There exists a constant $d_{0}>0$ such that
#$\#([x,x+x^{\theta}]\cap\mathcal{P})\geq\frac{d_{0}x^{\theta}}{\log x}$
for sufficiently large $x>0$.
Let $p_{1}$ be a sufficiently large prime number. By Theorem 2.1 with $x=p_{1}^{3}$, we find $p_{2}\in\mathcal{P}$ such that
$p_{1}^{3}\leq p_{2}\leq p_{1}^{3}+p_{1}^{3\theta}.$
Since $23\theta=63/40<$ 2 and $p_{1}$ is sufficiently large, we see that
$p_{1}^{3}+p_{1}^{3\theta}<p_{1}^{3}+3p_{1}^{2}+3p_{1}=(p_{1}+1)^{3}-1.$
Similarly, by Theorem 2.1 with $x=p_{2}^{3}$, we find $p_{3}\in\mathcal{P}$ such that
$p_{2}^{3}\leq p_{3}\leq p_{2}^{3}+p_{2}^{3\theta}<(p_{2}+1)^{3}-1.$
By iterating this argument, we find a sequence $(p_{k})_{k=1}^{\infty}$ of prime numbers such that
(2.1) $p_{k}^{3}\leq p_{k+1}\leq p_{k}^{3}+p_{k}^{3\theta}<(p_{k}+1)^{3}-1$
for every $k\in\mathbb{N}$. This leads to
$p_{1}^{1/3^{1}}\leq p_{2}^{1/3^{2}}\leq p_{3}^{1/3^{3}}\leq\cdots<(p_{3}+1)^{1/3^{3}}<(p_{2}+1)^{1/3^{2}}<(p_{1}+1)^{1/3^{1}},$
and hence $\lim_{k\to\infty}p_{k}^{1/3^{k}}\eqqcolon w$ and $\lim_{k\to\infty}(p_{k}+1)^{1/3^{k}}\eqqcolon w^{\prime}$ exist [^3] Actually, $w=w^{\prime}$ is true, but we do not use this formula.. Therefore, for every $k\in\mathbb{N}$, we have $p_{k}\leq w^{3^{k}}\leq w^{\prime 3^{k}}<p_{k}+1$, which implies that $p_{k}=\lfloor w^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$.
We will further give a proof of the following result for sharing the main idea.
Theorem 2.2.
The constant $w$ is irrational.
Remark 2.3.
We actually want to discuss the arithmetic properties of $\xi_{3}(=\min\mathcal{W}(3))$. There is a gap between $w$ and $\xi_{3}$. However, this gap is not essential to prove the irrationality (and transcendency). We can apply a similar way of $w$ to $\xi_{3}$ (see Lemmas 3.1 and 3.2).
To prove Theorem 2.2, we estimate the least distance between $w^{3^{k}}$ and prime numbers.
Lemma 2.4.
There exists $\gamma>0$ such that for every sufficiently large $k$
$|w^{3^{k}}-p_{k}|\leq e^{-3^{k}\gamma}.$
Proof.
Since $p_{k+1}=\lfloor w^{3^{k+1}}\rfloor$ for every $k\in\mathbb{N}$, (2.1) leads to
$p_{k}^{3}\leq p_{k+1}\leq w^{3^{k+1}}\leq p_{k}^{3}+p_{k}^{3\theta}+1\leq p_{k}^{3}(1+2p_{k}^{3(\theta-1)})$
for every sufficiently large $k\in\mathbb{N}$. Therefore,
$p_{k}\leq w^{3^{k}}\leq p_{k}(1+2p_{k}^{3(\theta-1)})^{1/3}.$
By applying the Maclaurin expansion, $(1+x)^{1/3}=1+O(x)$ holds if $|x|\leq 1$. Thus, since $\theta-1=-19/40<0$, for every sufficiently large $k$, we have
$p_{k}\leq w^{3^{k}}\leq p_{k}\left(1+O(p_{k}^{3(\theta-1)})\right)=p_{k}+O(p_{k}^{3(\theta-1)+1}),$
and hence there exists $K_{0}>0$ such that for every sufficiently large $k$
$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}.$
Combining $03\theta-2=63/40-2<$ 0 and the most left-hand side of (2.1), we obtain
$|w^{3^{k}}-p_{k}|\leq K_{0}p_{k}^{3\theta-2}\leq K_{0}p_{k-1}^{3(3\theta-2)}\leq\cdots\leq K_{0}p_{1}^{3^{k-1}(3\theta-2)},$
which completes the proof of Lemma 2.4. ∎
Lemma 2.4 means that $w^{3^{k}}$ is very close to some prime number. On the other hand, the powers of a rational number are at a distance from integers. Indeed, in 1957, Mahler [9] proved the following result. Here, for every $x\in\mathbb{R}$, we define
$\|x\|=\min\{|x-m|\colon m\in\mathbb{Z}\}.$
Theorem 2.5.
Let $r/s$ be a rational number greater than $1$ and which is not an integer. Let $\epsilon$ be a positive real number. Then, there exists an integer $n_{0}=n_{0}(r/s,\epsilon)$ such that
$\|(r/s)^{n}\|>e^{-\epsilon n}$
for every integer $n\geq n_{0}$.
Proof of Theorem 2.2.
We first assume that $w$ is an integer. Then
$p_{2}=\lfloor{w}^{9}\rfloor=(w^{3})^{3}=\lfloor w^{3}\rfloor^{3}=p_{1}^{3},$
which is a contradiction since $p_{2}$ is a prime number. Therefore, $w$ is not an integer. We secondly assume that $w$ is a rational number. Since $w$ is not an integer, there exist co-prime integers $r$ and $s$ with $r>s\geq 2$ such that $w=r/s$. By Lemma 2.4, there exists $\gamma>0$ such that for every sufficiently large $k$
$\|w^{3^{k}}\|\leq|w^{3^{k}}-p_{k}|\leq e^{-3^{k}\gamma}.$
On the other hand, by Theorem 2.5 with $n\coloneqq 3^{k}$ and $\epsilon\coloneqq\gamma$, for every sufficiently large $k$
$\|w^{3^{k}}\|>e^{-3^{k}\gamma}.$
This is a contradiction. Therefore, $w$ is irrational. ∎
Through the discussion above, we can share the main idea of the proof of Theorem 1.3. We will compare the following two phenomena:
  • (i)
    powers of the smallest PRC are very close to some prime number;
  • (ii)
    powers of almost all real algebraic numbers are at a distance from integers.
For (ii), we need a generalization of Theorem 2.5 to real algebraic numbers. Corvaja and Zannier [4] gave this generalization by applying the $p$-adic Schmidt subspace theorem. Dubickas used their result and showed the following [6, Lemma 6] to investigate a sequence $(x_{n})_{n=0}^{\infty}$ defined in (2.2).
Theorem 2.6.
Let $\alpha>1$ be an algebraic number, and $q$ be a positive integer. Suppose that $0<s_{1}<s_{2}<\cdots$ is a sequence of positive integers. Then, either for some $m\geq 0$ the number $\alpha^{s_{m}}$ is a Pisot number or for each $\epsilon>0$ there exists a positive integer $k_{0}=k_{0}(\epsilon)$ such that
$\|q\alpha^{s_{k}}\|>e^{-\epsilon s_{k}}$
for every $k\geq k_{0}$.
Let $P(x)$ be a polynomial of degree $d\geq 2$ with rational coefficients satisfying its leading coefficient is positive. Wagner and Ziegler [14] studied a sequence $(x_{n})_{n=0}^{\infty}$ satisfying
(2.2) $x_{n+1}=P(x_{n})\quad\text{for $n=0,1,2\ldots$, \quad and $x_{n}\to\infty$ as $n\to\infty$.}$
They showed that $\lim_{n\to\infty}x_{n}^{1/d^{n}}=\alpha>1$ exists and it is either irrational or an integer. Further, Dubickas [6] improved their result to the transcendency of $\alpha$ by applying Theorem 2.6. Their methods are very useful not only for the recursive equations (2.2) but also for recursive inequalities like (2.1). In addition, Dubickas presented the following lemma [6, Lemma 8].
Lemma 2.7.
Let $\beta$ be a Pisot number of degree $\ell\geq 2$ with conjugates $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ labelled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$. Then, there is a constant $\lambda>0$ depending only on $\beta$ such that
$|\beta_{2}^{n}+\cdots+\beta_{\ell}^{n}|\geq|\beta_{2}|^{n}n^{-\lambda}$
for each sufficiently large integer $n$.
We will apply Lemma 2.7 to the cases in which a Pisot number appears.
3. Lemmas
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3. By [13, Theorem 1.2], the set $\mathcal{W}(c_{k})$ is non-empty, closed, and lower bounded. Therefore, $\xi\coloneqq\min\mathcal{W}(c_{k})$ exists. We may assume that $(c_{k})_{k=1}^{\infty}$ is bounded, if not, $\xi$ is transcendental by [13, Theorem 1.3]. Thus setting $B=\sup_{k\in\mathbb{N}}c_{k}$, we have $B<\infty$. Let $\mathcal{I}_{b}=\{k\in\mathbb{N}\colon c_{k+1}=b\}$ which is an infinite set from (3). Let $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for all $k\in\mathbb{N}$. Let $\theta=21/40$. The following Lemmas 3.1 and 3.2 can be seen in [13, Lemma 4.5 and Lemma 5.1], but we give proofs for keeping the readability.
Lemma 3.1.
For all $k\in\mathbb{N}$, we have
(3.1) $p_{k}^{c_{k+1}}\leq p_{k+1}<(p_{k}+1)^{c_{k+1}}-1.$
Proof.
Take any $k\in\mathbb{N}$. By the definition of $\xi$, we see that $p_{k}\leq\xi^{C_{k}}<p_{k}+1$, which implies that $p_{k}^{c_{k+1}}\leq\xi^{C_{k+1}}<(p_{k}+1)^{c_{k+1}}$. We recall that $c_{k+1}$ is a positive integer greater than or equal to $2$, and hence $p_{k}^{c_{k+1}}\leq\lfloor\xi^{C_{k+1}}\rfloor<(p_{k}+1)^{c_{k+1}}$. It is not true that $p_{k+1}=(p_{k}+1)^{c_{k+1}}-1$ since $p_{k+1}$ is a prime number. Therefore, we obtain (3.1). ∎
Lemma 3.2.
There exists $k_{0}=k_{0}(\xi)>0$ such that for all $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$, we have
(3.2) $p_{k}^{c_{k+1}}\leq p_{k+1}\leq p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}.$
To prove Lemma 3.2, we apply the following result given by Matomäki (see [11, Lemma 9] and [10]).
Theorem 3.3.
There exist positive constants $d_{1}<1$ and $D$ such that, for every sufficiently large $x$ and every $\gamma\in[1/2,1]$, the interval $[x,2x]$ contains at most $Dx^{2/3-\gamma}$ disjoint intervals $[n,n+n^{\gamma}]$ for which
#$\#([n,n+n^{\gamma}]\cap\mathcal{P})\leq d_{1}\frac{n^{\gamma}}{\log n}.$
By applying this theorem, Matomäki [11] constructed an element in $\mathcal{W}(2)$. It is truly astonishing because Mills’ construction tells us that we need the existence of a prime number in each interval $[n,n+n^{1/2}]$ to concrete an element in $\mathcal{W}(2)$. This problem is strongly connected with Legendre’s conjecture which asserts that for every $n$ there exists a prime number between $n^{2}$ and $(n+1)^{2}$. It is unsolved and believed to be extremely difficult. However, Theorem 3.3 with $\gamma=1/2$ states that there exists a prime number $p\in[n,n+n^{1/2}]$ except for very few intervals $[n,n+n^{1/2}]$. Therefore, Matomäki discovered that we can construct a PRC if the exceptional intervals are very few. We apply Theorem 3.3 to prove the following lemma.
Lemma 3.4.
Let $c$ be a real number in the interval $[2,B]$, and $\epsilon$ be a sufficiently small positive real number. Let $X_{0}$ be a sufficiently large positive parameter depending only on $B$ and $\epsilon$. Let $X$ and $\eta$ be real numbers satisfying $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$. Suppose that
(3.3) #$\#([X,X+X^{\eta}]\cap\mathcal{P})\geq\frac{d_{2}X^{\eta}}{\log X}$
for some constant $d_{2}>0$. Then there exists $q\in[X,X+X^{\eta}]\cap\mathcal{P}$ such that
(3.4) #$\#([q^{c},q^{c}+q^{c-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c-1}}{\log q^{c}}.$
Proof.
Let $R=[X,X+X^{\eta}]\cap\mathcal{P}$. Each interval $[q^{c},q^{c}+q^{c-1}]$ $(q\in R)$ is completely contained in $[X^{c},2X^{c}]$. Indeed, we have
$X^{c}$ $\leq q^{c}<q^{c}+q^{c-1}=q^{c}(1+q^{-1})\leq(X+X^{\eta})^{c}(1+X^{-1})$
$\leq X^{c}(1+X^{\eta-1})^{B}(1+X^{-1})\leq 2X^{c}$
since $\eta\leq 1-\epsilon$ and $X$ is sufficiently large. Setting $\gamma=(c-1)/c$, we have $\gamma\in[1/2,1)$. By $\eta\geq 1/2$ and $c\geq 2$, the following inequalities are true:
(3.5) $\frac{d_{2}X^{\eta}}{\log X}-DX^{c(2/3-\gamma)}$ $=\frac{d_{2}X^{\eta}}{\log X}-DX^{1-c/3}=\frac{X^{\eta}}{\log X}\left(d_{2}-DX^{1-\eta-c/3}\log X\right)$
$\geq\frac{{X^{\eta}}}{\log X}\left(d_{2}-DX^{-1/6}\log X\right)>0,$
since $X\geq X_{0}$ and $X_{0}$ is sufficiently large. Then by combining (3.3), (3.5), and Theorem 3.3 with $x\coloneqq X^{c}$ and $\gamma\coloneqq\gamma$, there exists a prime number $q\in R$ satisfying (3.4). ∎
Proof of Lemma 3.2.
Assume that for every $k_{0}>0$ there exists $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$ such that
(3.6) $p_{k+1}\notin[p_{k}^{c_{k+1}},p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}].$
We will find $w\in\mathcal{W}(c_{k})$ such that $w<\xi$ by Matomäki’s construction, which will imply a contradiction to the minimality of $\xi$.
We now take $k_{0}$ as a sufficiently large positive parameter. By the assumption, we find an integer $k$ in $\mathcal{I}_{b}\cap[k_{0},\infty)$ satisfying (3.6). By Lemma 3.1, it is false that $p_{k}^{c_{k+1}}>p_{k+1}$, and hence we have
(3.7) $p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}<p_{k+1}.$
Then, we set $q_{k}=p_{k}$. Theorem 2.1 with $x\coloneqq q_{k}^{c_{k+1}}$ implies that
#$\#([q_{k}^{c_{k+1}},q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}]\cap\mathcal{P})\geq\frac{d_{0}q_{k}^{\theta c_{k+1}}}{\log q_{k}^{c_{k+1}}}.$
By Lemma 3.4 with $X\coloneqq q_{k}^{c_{k+1}}$, $\eta\coloneqq\theta$, and $c\coloneqq c_{k+2}$, there exists $q_{k+1}\in\mathcal{P}$ satisfying
(3.8) $q_{k}^{c_{k+1}}\leq q_{k+1}\leq q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}(\leq q_{k}^{c_{k+1}}+q_{k}^{c_{k+1}-1})$
(3.9) #$\#([q_{k+1}^{c_{k+2}},q_{k+1}^{c_{k+2}}+q_{k+1}^{c_{k+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c_{k+2}-1}}{\log q^{c_{k+2}}}.$
Let us construct a sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying the following properties: for every $m\geq k$, we have
(3.10) $q_{m}^{c_{m+1}}\leq q_{m+1}\leq q_{m}^{c_{m+1}}+q_{m}^{c_{m+1}-1},$
(3.11) #$\#([q_{m+1}^{c_{m+2}},q_{m+1}^{c_{m+2}}+q_{m+1}^{c_{m+2}-1}]\cap\mathcal{P})\geq\frac{d_{3}q_{m+1}^{c_{m+2}-1}}{\log q_{m+1}^{c_{m+2}}},$
where $d_{3}$ is a small positive constant. In the case $m=k$, both (3.10) and (3.11) are true by (3.8) and (3.9). We assume that there exists a sequence $(q_{m})_{m=k}^{M}$ of prime numbers satisfying (3.10) and (3.11) for every $m=k,\ldots,M-1$, where $M\geq k+1$. Then by applying Lemma 3.4 with $X\coloneqq q_{M}^{c_{M+1}}$, $\eta\coloneqq(c_{M+1}-1)/c_{M+1}$, and $c\coloneqq c_{M+2}$, there exists a prime number $q_{M+1}$ satisfying (3.10) and (3.11) with $m=M$. By induction, we obtain an infinite sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying (3.10) and (3.11).
Furthermore, we set $q_{m}=p_{m}$ for every $m=1,2,\ldots,k-1$. By Lemma 3.1 and (3.10), we obtain
$q_{m}^{c_{m+1}}\leq q_{m+1}<(q_{m}+1)^{c_{m+1}}-1$
for every $m\in\mathbb{N}$. Therefore, we have
(3.12) $q_{m}^{1/C_{m}}\leq q_{m+1}^{1/C_{m+1}}<(q_{m+1}+1)^{1/C_{m+1}}<(q_{m}+1)^{1/C_{m}}$
for every $m\in\mathbb{N}$, and hence $\lim_{m\to\infty}q_{m}^{1/C_{m}}=w$ and $\lim_{m\to\infty}(q_{m}+1)^{1/C_{m}}=w^{\prime}$ exist. In addition, $w\in\mathcal{W}(c_{k})$ holds by (3.12). On the other hand, combining (3.7) and (3.8),
$\lfloor w^{C_{k+1}}\rfloor=q_{k+1}\leq p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}}<p_{k+1}=\lfloor\xi^{C_{k+1}}\rfloor.$
This is a contradiction since $\xi$ is the smallest element in $\mathcal{W}(c_{k})$. ∎
Lemma 3.5.
There exist $k_{1}=k_{1}(\xi)>0$ and $K_{0}>0$ such that for every integer $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$, we have
(3.13) $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}},$
where $\theta_{b}=1-\theta-1/b>0$. Further, there exists $\gamma>0$ such that for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$
(3.14) $|\xi^{C_{k}}-p_{k}|\leq e^{-\gamma C_{k}}.$
Proof.
Let $k_{0}(\xi)$ be as in Lemma 3.2, and let $k_{1}=k_{1}(\xi)>k_{0}(\xi)$ be a sufficiently large parameter. Since $p_{k}=\lfloor\xi^{C_{k}}\rfloor$ for every $k\in\mathbb{N}$, by Lemma 3.2, each $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$ satisfies
(3.15) $p_{k}^{c_{k+1}}\leq p_{k+1}\leq\xi^{C_{k+1}}\leq p_{k+1}+1\leq p_{k}^{c_{k+1}}+2p_{k}^{\theta c_{k+1}}.$
Therefore, for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$, we have
(3.16) $p_{k}\leq\xi^{C_{k}}\leq(p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}})^{1/c_{k+1}}=p_{n}(1+p_{k}^{c_{k+1}(\theta-1)})^{1/c_{k+1}}=p_{k}+O(p_{k}^{c_{k+1}(\theta-1)+1}).$
We note that $c_{k+1}(\theta-1)+1=c_{k+1}(\theta-1+\frac{1}{b})=-c_{k+1}\theta_{b}<0$ since $c_{k+1}=b$ and $b\geq 3$. Therefore, we get (3.13). Furthermore, Lemma 3.1 leads to
(3.17) $p_{k}^{c_{k+1}}\geq p_{k-1}^{c_{k+1}c_{k}}\geq p_{k-2}^{c_{k+1}c_{k}c_{k-1}}\geq\cdots\geq p_{1}^{C_{k+1}/c_{1}}.$
Combining (3.13) and (3.17), we obtain (3.14). ∎
Remark 3.6.
We can observe the main difficulties of Question 1.4 in (3.16). If $c_{1}=c_{2}=\cdots=2$, then roughly speaking, (3.16) should be
$p_{k}\leq\xi^{2^{k}}<p_{k}+p_{k}^{2(1/2-1)+1}=p_{k}+1.$
It does not suffice to obtain (3.14). Further, by Matomäki’s construction, we do not know which specific prime number $q$ satisfies (3.4). This is also one of the main difficulties.
4. Proof of Theorem 1.3
Lemma 4.1.
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3. Let $\xi$ be the smallest element in $\mathcal{W}(c_{k})$. Then the following properties are true:
  • if $b\geq 5$, then $\xi$ is transcendental;
  • if $b=4$, then either $\xi$ is transcendental, or there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$;
  • if $b=3$, then either $\xi$ is transcendental, or there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$ or $3$.
Proof.
Let $k_{1}$ be as in Lemma 3.5. Assume that $\xi$ is a real algebraic number. By the definition of $\mathcal{W}(c_{k})$, $\xi>1$ holds. By Lemma 3.5, there exists $\gamma>0$ such that for every $k\in\mathcal{I}_{b}\cap[k_{1},\infty)$
$|\xi^{C_{k}}-p_{k}|\leq e^{-\gamma C_{k}}.$
Therefore, Theorem 2.6 with $\alpha\coloneqq\xi$, $q\coloneqq 1$, $s_{k}\coloneqq C_{k}$, and $\epsilon\coloneqq\gamma$ implies that $\xi^{C_{m}}$ is a Pisot number for some $m\in\mathbb{N}$. Let $\beta=\xi^{C_{m}}$, and let $\ell$ be the degree of $\beta$. We now show the following claim, where $\theta_{b}$ is defined in Lemma 3.5.
Claim. We have $12\leq\ell\leq(b\theta_{b})^{-1}+$ 1.
If $\ell=1$, then $\beta(=\xi^{C_{m}})$ is a positive integer. Therefore,
$p_{m+1}=\lfloor\xi^{C_{m+1}}\rfloor=(\xi^{C_{m}})^{c_{m+1}}=p_{m}^{c_{m+1}},$
which is a contradiction since $p_{m+1}$ is a prime number and $c_{m+1}\geq 2$. Therefore $\ell\geq 2$.
We next assume that $\ell>(b\theta_{b})^{-1}+1$. Let $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ be the conjugates over $\mathbb{Q}$ of $\beta$ labeled so that $\beta_{1}>1>|\beta_{2}|\geq\cdots\geq|\beta_{\ell}|$. For every $k\geq m$, since $C_{k}/C_{m}$ is a positive integer, we obtain
$\beta_{1}^{C_{k}/C_{m}}+\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\in\mathbb{Z}.$
Let $k$ be a sufficiently large integer in $\mathcal{I}_{b}\cap[m,\infty)$. By Lemma 3.5,
(4.1) $\|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\|=\|\beta_{1}^{C_{k}/C_{m}}\|=\|\xi^{C_{k}}\|\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}.$
By (3.15), we recall that $\xi^{C_{k+1}}\leq p_{k}^{c_{k+1}}+2p_{k}^{\theta c_{k+1}}$, to obtain
(4.2) $\xi^{C_{k+1}}\leq p_{k}^{c_{k+1}}(1+2p_{k}^{(\theta-1)c_{k+1}})\leq 2p_{k}^{c_{k+1}}.$
Therefore, combining (4.1), (4.2), and $c_{k+1}=b$, we have
$\|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\|\leq K_{1}\xi^{-C_{k+1}\theta_{b}}=K_{1}\beta_{1}^{-b\theta_{b}C_{k}/C_{m}}$
for some constant $K_{1}>0$. By Lemma 2.7 with $n\coloneqq C_{k}/C_{m}$, there exists $\lambda=\lambda(\beta)>0$ such that
$\left\|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\right\|=\left|\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}\right|\geq|\beta_{2}|^{C_{k}/C_{m}}(C_{k}/C_{m})^{-\lambda}$
since $k$ is sufficiently large. By taking $k\to\infty$ on $k\in\mathcal{I}_{b}$, we conclude that $|\beta_{2}|\leq\beta_{1}^{-b\theta_{b}}$. Since $|\beta_{2}|\geq|\beta_{3}|\geq\cdots\geq|\beta_{l}|$ and $\ell>(b\theta_{b})^{-1}+1$, we have $|\beta_{2}\cdots\beta_{\ell}|\leq\beta_{1}^{-(\ell-1)b\theta_{b}}<\beta_{1}^{-1}$. This is a contradiction to $|\beta_{1}\cdots\beta_{\ell}|\geq 1$. Thus, we obtain the claim.
Case ($b\geq 5$): It follows that
$b\theta_{b}\geq\frac{19\cdot 5}{40}-1\geq\frac{19}{8}-1=\frac{11}{8}\quad\therefore\quad(b\theta_{b})^{-1}+1\leq\frac{8}{11}+1<2,$
which is a contradiction to the claim. Therefore, $\xi$ is transcendental.
Case ($b=4$): By simple calculation,
$b\theta_{b}=\frac{19\cdot 4}{40}-1=\frac{19}{10}-1=\frac{9}{10}\quad\therefore\quad(b\theta_{b})^{-1}+1=\frac{10}{9}+1<3.$
Case ($b=3$): We also observe that
$b\theta_{b}=\frac{19\cdot 3}{40}-1=\frac{57}{40}-1=\frac{17}{40}\quad\therefore\quad\ (b\theta_{b})^{-1}+1=\frac{40}{17}+1<4.$
This completes the proof. ∎
Let us discuss the case when $\xi^{C_{m}}$ is a Pisot number of degree $2$ for some $m\in\mathbb{N}$.
Lemma 4.2.
Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. Suppose that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number. Let $\beta=\xi^{C_{m}}$ and $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ be the conjugates over $\mathbb{Q}$ of $\beta$. We set
$t_{k}=\beta_{1}^{C_{k}/C_{m}}+\cdots+\beta_{\ell}^{C_{k}/C_{m}}\in\mathbb{Z}$
for every $k\geq m$. Then we have $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$ for every sufficiently large $k\in\mathbb{N}$.
Proof.
Let $\delta_{k}=\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}$ for every $k\geq m$. Let $k$ be a sufficiently large integer in $\mathcal{I}_{b}\cap[m,\infty)$. Since $t_{k}$ and $\beta_{1}^{C_{k}/C_{m}}$ are real numbers, $\delta_{k}$ is also a real number. Further, we recall that $|\beta_{j}|<1$ for every $2\leq j\leq\ell$, which implies that $\delta_{k}\in(-1/2,1/2)$ since $k$ is sufficiently large. If $-1/2<\delta_{k}\leq 0$, then $\xi^{C_{k}}=\beta_{1}^{C_{k}/C_{m}}=t_{k}-\delta_{k}$, and hence $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$. If $20<\delta_{k}<1/$ 2, then by (3.13), $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}$ holds for some constant $K_{0}>0$. Therefore,
$\delta_{k}\leq t_{k}-p_{k}\leq\delta_{k}+K_{0}p_{k}^{-c_{k+1}\theta_{b}}.$
Since $t_{k}-p_{k}\in\mathbb{Z}$ and $-c_{k+1}\theta_{b}<0$, this is a contradiction by choosing $k$ large enough. ∎
Lemma 4.3.
Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. If $b\in\{3,4\}$, then there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$.
Proof.
Assume that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$. Let $\beta=\xi^{C_{m}}$, and let $\beta_{1}=\beta,\beta_{2}$ be the conjugates of $\beta$. Let $t_{k}$ be as in Lemma 4.2.
Case ($b=3$): For every $k\in\mathcal{I}_{3}$, we set $x_{1}=\beta_{1}^{C_{k}/C_{m}}$ and $x_{2}=\beta_{2}^{C_{k}/C_{m}}$. Then we obtain $c_{k+1}=3$ and
$t_{k}^{3}=(x_{1}+x_{2})^{3}=x_{1}^{3}+3x_{1}^{2}x_{2}+3x_{1}x_{2}^{2}+x_{2}^{3}=t_{k+1}+3x_{1}x_{2}t_{k}.$
By applying Lemma 4.2, for sufficiently large $k\in\mathcal{I}_{3}$, we have $p_{k}^{3}=p_{k+1}+3x_{1}x_{2}p_{k}$. By $x_{1}x_{2}\in\mathbb{Z}$, we see that $p_{k+1}$ is divisible by $p_{k}$. This is a contradiction since $p_{k}$ and $p_{k+1}$ are distinct prime numbers.
Case ($b=4$): Let $k$ be a sufficiently large integer in $\mathcal{I}_{4}\cap[m,\infty)$. Since $\mathcal{I}_{4}$ is infinite, we may assume that $C_{k}/C_{m}(\eqqcolon d_{k})$ is a positive even integer. Setting $x_{1}=\beta_{1}^{d_{k}}$ and $x_{2}=\beta_{2}^{d_{k}}$, we have
$\beta_{1}^{d_{k}}+\beta_{2}^{d_{k}}=p_{k}=\lfloor\beta_{1}^{d_{k}}\rfloor\leq\beta_{1}^{d_{k}}.$
Then $\beta_{2}^{d_{k}}\leq 0$. This is a contradiction since $\beta_{2}\in\mathbb{R}$ and $d_{k}$ is a positive even number. ∎
Combining Lemmas 4.1 and 4.3 completes the proof of Theorem 1.3.
Remark 4.4.
To prove the transcendency of Mills’ constant, it remains to show that there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $3$ when $b=3$. In this remark, we describe the difficulty of this case.
Suppose $b=3$, and we assume that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $3$. Let $\beta=\xi^{C_{m}}$ and $\beta_{1}=\beta,\beta_{2},\beta_{3}$ be the conjugates of $\beta$. Similarly with the proof of Lemma 4.3, we obtain
(4.3) $p_{k+1}=p_{k}^{3}-3b_{k}p_{k}+3e_{k}$
for every sufficiently large $k\in\mathcal{I}_{3}$, where
$b_{k}\coloneqq(\beta_{1}\beta_{2})^{C_{k}/C_{m}}+(\beta_{2}\beta_{3})^{C_{k}/C_{m}}+(\beta_{3}\beta_{1})^{C_{k}/C_{m}},\quad e_{k}\coloneqq(\beta_{1}\beta_{2}\beta_{3})^{C_{k}/C_{m}}.$
Therefore, we have to investigate properties of $b_{k}$ and $e_{k}$. The author does not have any good ideas on how to treat $b_{k}$ and $e_{k}$ simultaneously. Indeed, the method of the proof of Lemma 4.3 does not work.
5. Further discussions
We will give several assumptions leading to the transcendency of $\xi_{3}$.
Proposition 5.1.
Assume that $\lfloor\xi_{3}^{3^{k}}\rfloor\not\equiv\lfloor\xi_{3}^{3^{k+1}}\rfloor\mod 3$ for infinitely many positive integers $k$. Then $\xi_{3}$ is transcendental.
Proof.
Suppose that $\xi_{3}$ is algebraic. Then by Theorem 1.2, there exists $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. Let $p_{k}=\lfloor\xi_{3}^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$. By Fermat’s little theorem, $p_{k}^{3}\equiv p_{k}\mod 3$. Thus, by Remark 4.4, we have $p_{k+1}\equiv p_{k}\mod 3$ for every sufficiently large $k\geq m$. This is a contradiction. ∎
A Mills prime is a prime number of the form $\lfloor\xi_{3}^{3^{k}}\rfloor$. By Proposition 5.1, if there are infinitely many Mills primes of the forms $13n+$ 1 and $23n+$ 2 respectively, then $\xi_{3}$ is transcendental. However, it seems to be impossible to classify Mills primes modulo $3$.
Proposition 5.2.
Assume that there exist $\theta^{\prime}\in(0,1/2)$ and $x_{0}>0$ such that for every $x\geq x_{0}$ we can find a prime number $p\in[x,x+x^{\theta^{\prime}}]$. Then $\xi_{3}$ is transcendental.
Proof.
Suppose that $\xi_{3}$ is algebraic. Then by Theorem 1.2, there exists $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. We set $\theta_{b}^{\prime}=\theta_{3}^{\prime}=1-\theta^{\prime}-1/3$. In a similar manner with Lemma 3.5 and the claim in the proof of Lemma 4.1, we have
$3=\deg(\xi_{3}^{3^{m}})\leq(3\theta_{3}^{\prime})^{-1}+1=\frac{1}{2-3\theta^{\prime}}+1<\frac{1}{2-3/2}+1=3,$
which is a contradiction. ∎
Let $p_{n}$ be the $n$th prime number for every $n\in\mathbb{N}$. Cramér [5] conjectured that $p_{n+1}-p_{n}=O((\log p_{n})^{2})$. If this conjecture is true, we can take $\theta^{\prime}$ in Proposition 5.2 as an arbitrarily small positive constant. Therefore, assuming Cramér’s conjecture, $\xi_{3}$ is transcendental.
Acknowledgement
The author would like to thank Professor Shigeki Akiyama for daily attractive discussions on mathematics. The author is grateful to Professor Hajime Kaneko for giving useful ideas on the cases $b=3$ and $4$. The author is also grateful to Professor Wataru Takeda for finding errors and mistakes. The author was supported by JSPS KAKENHI Grant Numbers JP22J00025 and JP22KJ0375.